0.5 mL of a 0.04 M CoCl2 solution is combined with 1.5 mL deionized water. What is the molarity of the solution?
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 0.04 M

V1 = 0.5 mL

V2 = 1.5 mL + 0.5 mL = 2 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (0.04*0.5)/2

M2 = 0.01 M

Answer: 0.01 M